14.37 Using the data in Problem 14.36, construct a 95% confidence interval for the regression slope.Rating 16 Rating 15 18 12 13 Increase ($) 2,400 2,500 1,800 1,700 15 17 16 Increase ($) 2,000 2,600 2,000 1,900
Answer #1

SUMMARY OUTPUT Regression Statistics Multiple R 0.588422802 R Square 0.346241394 Adjusted R Square 0.237281626 Standard ErrorUsing excel the data analysis are Slope by = 100.9091 Standard error for the slope Se(b1)=56.6075 Df=n-2=8-2 =6 Level of sign